OpenStudy (anonymous):

What is the derivative of 3x^2 + 5x +10? Can someone help explain please?

OpenStudy (anonymous):

6x+5

OpenStudy (anonymous):

Can you help explain that please?

OpenStudy (anonymous):

To step through, we can look at the function 3x^2 + 5x + 10 as three functions added together

OpenStudy (anonymous):

use sum rule n power rule

OpenStudy (anonymous):

so consider, the derivative of 10...10 is a constant

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

d/d(3x^2)=6x d/dx 5x=5 d/dx(10)=0

OpenStudy (anonymous):

looks like you want derivative of the function with respect to x...therefore constants become 0...right check out uzma's work

OpenStudy (anonymous):

so, as uzma found a first response: \[6x + 5\]

OpenStudy (anonymous):

Thanks! any more questions?